OpenStudy (anonymous):

Train A leaves New York at 7:00am traveling to Boston at 80mph. Train B leaves Boston at 7:45am traveling to New York at 70mph on a parallel track. If the distance between New York and Boston along the tracks is 210 miles, at what time will the two trains pass each other?

OpenStudy (anonymous):

HaHa We meet Again

hero (hero):

I can do this on twiddla

OpenStudy (anonymous):

PASS each other

OpenStudy (anonymous):

I can do this here

OpenStudy (anonymous):

Just a minute

hero (hero):

I can too, but it's better on twiddla.

hero (hero):

I can draw pictures

hero (hero):

and charts

OpenStudy (anonymous):

http://www.twiddla.com/588187

OpenStudy (anonymous):

Your Question is Wrong

OpenStudy (anonymous):

The Faster train starts at 7 and slower one at 7:45 How will they ever PASS

OpenStudy (anonymous):

they are from two separate cities

OpenStudy (anonymous):

going to one another

OpenStudy (anonymous):

My Bad I should improve reading Sorry

hero (hero):

I get 1.05 hrs

OpenStudy (anonymous):

Okay So See my Creative Mind 0 210 ||___________________________________|____________________________________________|| Boston x N.Y Now we have it that they meet at x distance form Boston Now we have two equations \[\frac{210 - x}{70} = t + \frac{3}{4}\] \[\frac{x}{80} = t\]

hero (hero):

So, at just over 8:00, then will pass

hero (hero):

How'd you end up with two variables?

hero (hero):

I only needed one equation

OpenStudy (anonymous):

Well it is 210-x for one train and x for other

OpenStudy (anonymous):

but the difference in time is 3/4 hours

hero (hero):

Actually, the exact time is 8:03

hero (hero):

At 8:03, they will pass

hero (hero):

GEEZ, they need to fix that...........

OpenStudy (anonymous):

So x/80 is the time taken by faster one then 210-x is for the slower one ----- 70 Now because they have 3/4 hr difference t = x/80 t - 3/4 = 210-x ------- 70

OpenStudy (anonymous):

ah had a typo before it's - not + 3/4

hero (hero):

I posted my solution on Twiddla...it took less than two minutes total time

OpenStudy (anonymous):

Hero think about it in one hour the fast one shall travel 80 miles then 70 miles one should travel 70./4 how could they meet if 210 is the distance

OpenStudy (anonymous):

i mean in 1 hour

OpenStudy (anonymous):

I agree

OpenStudy (anonymous):

One hour is too short

OpenStudy (anonymous):

It should be around 1 hr and 40 minutes something like that

hero (hero):

Yeah, I recalculated and got 1hr 36 minutes

hero (hero):

So 8:36

OpenStudy (anonymous):

using the same formula?

hero (hero):

No, using a different one

OpenStudy (anonymous):

what is the proper equation?

hero (hero):

Wait, let me verify all of this..

OpenStudy (anonymous):

i am getting 1.7 hours at x=140

OpenStudy (anonymous):

and this should be quite right 1.75*80 = 140 and 70*1 =70

OpenStudy (anonymous):

so we have 1.75 hours from 7:00 that makes 8:45 correct time

OpenStudy (anonymous):

damn sure hafeez i am 100% correct

OpenStudy (anonymous):

Hero it took us more then 2 minutes lol

OpenStudy (anonymous):

Hey hit the good answer

hero (hero):

Yeah, actually...I had the right equation, I just subtracted when I should have added

hero (hero):

The equation I posted on Twiddla is right....

OpenStudy (anonymous):

lolz it is mathematical error with human emotions

hero (hero):

Just add 52.5 to 210 instead of subtract

hero (hero):

Ishaan, I can't clarify a mistake?

hero (hero):

I guess I'm not allowed to do that...

OpenStudy (anonymous):

Okay Dokey :D

hero (hero):

x = 1.75

OpenStudy (anonymous):

time =1.75 yeah right where is hafeez he is not here

hero (hero):

Of course not

hero (hero):

http://www.twiddla.com/588187

hero (hero):

I posted my solution here. You can see where I made my mistake